# Anonymous Authors    Carlos Güemes-Palau, Miquel Ferriol-Galmés, Jordi Paillisse Vilanova, Albert López-Brescó, Pere Barlet-Ros and Albert Cabellos-Aparicio

### v1

@@ I Introduction
Network modeling is crucial for reliable communication, traffic optimization, and network design. Traditional approaches, such as Discrete Event Simulation (DES), have been the standard for accurately modeling network behavior. Tools like NS-3 [1] and OMNeT++ [2] simulate every network event. They provide highly detailed reconstructions of network dynamics but at a high computational cost, which limits scalability as networks grow. Recent advancements in Machine Learning (ML) have introduced transformative possibilities for network modeling, exemplified by models such as MimicNet [3], DeepQueueNet [4], and the RouteNet [5] family. These innovations address the limitations of DES, enabling faster and more scalable network modeling at the price of a slight reduction in accuracy.
@@
Researchers often turn to network simulators to generate training datasets to mitigate data scarcity [3, 4, 5, 7]. Simulators offer a controlled and flexible environment for producing diverse and rare scenarios without the risks associated with live production networks. Conversely, generating simulated datasets is computationally expensive and often fails to capture the subtle nuances of real-world network dynamics. This arises from idealized assumptions in simulations and the proprietary nature of commercial network hardware. Consequently, ML models trained exclusively on simulated data exhibit reduced accuracy and reliability when deployed in production environments. This observation has been corroborated by other researchers [8] and by our findings in the evaluation (Table II).
This paper studies the viability of using simulated data to train ML models for real-world network behavior prediction. We propose a hybrid approach leveraging transfer learning [9, 10] to bridge the gap between simulation and real-world network environments. This consists in training and evaluating our approach by combining simulated network scenarios with real ones (Figure 1). We use RouteNet-Fermi [5], a state-of-the-art ML network model, as our reference model in our comparison. However, we believe that our findings are generalizable to other ML-based solutions. We then use our configurable testbed network to record the real network scenarios. In our evaluation, we increase the accuracy of the resulting model by up to 88% relative to the version that only uses real-world network data. We also find that our approach is extremely data efficient: limiting the training process to 10 recorded network scenarios, using fine-tuning results in a model with $37\%$ less error than when training from scratch.
@@ II-A Transfer Learning
    Reused and adjustable weights: The weights are transferred from the donor model but are allowed to update during training —and thus adapt to better fit the target.
@@ II-B RouteNet-Fermi
RouteNet-Fermi [5] is a state-of-the-art network performance model. Specifically, it uses a custom representation of the network along with a modified Message-Passing Neural Network architecture [13] that exploits the interactions between traffic flows and the underlying devices. As a result, it has proven to generate accurate predictions at a fraction of the computational cost of alternatives such as DES. Its design also makes it robust to infer under unseen topologies, including larger than those seen during the model’s training. It was also evaluated with both simulated and real-world network data.
@@ IV Methodology and experimental design
Our approach involves two main components: (1) training the RouteNet-Fermi [5] model on simulated data and (2) fine-tuning it with real-world network data. This methodology enables us to combine the broad generalization capabilities gained from simulation with real-world specificity.
@@ IV-B Manual transfer learning
The resulting testbed configurations are listed in Table II in the evaluation. We split the network into blocks rather than individual layers to align with RouteNet-Fermi’s architecture. Unlike traditional NNs like MLPs, which are sequential, RouteNet-Fermi operates more like an ensemble of smaller NNs that work in parallel. For instance, the MLPs in the encoding block process individual network elements independently. Grouping these layers into blocks provides a structured approach to fine-tuning while ensuring that dependencies between blocks are respected. Furthermore, the shallow depth of the RouteNet-Fermi’s internal NNs, with the deepest component being a 3-layered MLP, limits the benefit of fine-tuning individual layers.
Figure 2: Visual example of fine-tuning a RouteNet-Fermi [5] model, where the Encoding is frozen, the MPA is fine-tuned, and the Readout is re-trained.
@@ IV-C Automated transfer learning
  • Autofreeze [15]: This method consists of loading all donor weights and setting them as trainable. During training, blocks whose weight gradients fall under a threshold are frozen, allowing weights to be adjusted while minimizing computational costs.
@@
  • GTOT-Tuning [17]: A more advanced version of L2-SP meant for Graph Neural Networks (GNNs). Instead of comparing weights, it measures the differences between node embeddings after the MPA using the Masked Wasserstein Distance (MWD). The mask in the MWD allows it to incorporate relational information.
@@ V Evaluation
TABLE I: Number of network scenarios per source, traffic distribution, and partition.
@@
Table I presents the number of network scenarios available for our study, categorized by source (simulated or real), traffic distribution (Poisson, On/Off, or MAWI), and partition (training, validation, or evaluation). This partitioning ensures that the evaluation remains unbiased and robust. Notably, the simulated dataset is significantly larger than the real-world dataset, reflecting the practical challenge of limited availability and variability in real-world network data.
As part of the evaluation, we establish two baseline models: one trained exclusively on simulated scenarios and the other solely on real-world scenarios. We then use the simulated data only model for both manual and automated fine-tuning. Manual fine-tuning configurations and criteria for selecting these configurations are detailed in subsections IV-B. The automated fine-tuning approaches are listed in Section IV-C and have been implemented according to their original papers, including hyperparameter values, with necessary adaptations made to fit RouteNet-F’s architecture as needed. Our windowed implementation of the RouteNet-Fermi model is built upon their public repository [^1] https://github.com/BNN-UPC/RouteNet-Fermi, using TensorFlow 2.15. Model hyperparameters (e.g., embedding sizes) are borrowed directly from the original implementation. All models ran for a redundant number of epochs until the validation error stopped improving. Then, the model kept the checkpoint weights at which the error was minimized. Additional implementation details for this work are available in our public repository [^2] https://github.com/BNN-UPC/Papers/wiki/Bridging-the-Gap-Between-Simulated-and-Real-Network-Data-Using-Transfer-Learning.
@@ V-A Fine-tuning the network model
Table II presents the MAPE for all evaluated fine-tuning configurations. Models are tasked to predict the average packet delay in each flow-window pair. The manual configurations involve varying the treatment of weights (freezing, fine-tuning, or re-training) in the three main blocks of the RouteNet-Fermi model: Encoding, MPA, and Readout. We also show the performance of the models trained exclusively on simulated data and automated finetuning methods. The results are normalized against a model trained with the real-world network data only, without fine-tuning. That is, a value of 1 indicates the same accuracy as the baseline, while lower values indicate an improvement over the baseline.
@@ V-B Impact of real-world network data availability
The results show that with only five scenarios available, fine-tuning reduces the model’s MAPE from $19.01\%$ to $16.19\%$. Increasing the number of available scenarios to 10 further decreases the MAPE to $11.95\%$, a $37\%$ reduction. With 50 scenarios, the fine-tuned model achieves its largest improvement, lowering the MAPE to $9.95\%$, a $48\%$ decrease. This demonstrates the data efficiency enabled by fine-tuning. However, these results also reveal their limitations: beyond 125 scenarios, fine-tuning provides no additional benefit over training with real-world network data alone. While this threshold may vary between models, RouteNet-Fermi has demonstrated the ability to achieve high accuracy even with limited training data [5]. This reinforces the idea that fine-tuning is particularly valuable when real-world network data is scarce.
@@ VI-A Network modeling
State-of-the-art network models can be broadly categorized into two fields: DES and ML-based models. DES simulators such as OMNeT++[2] and NS3[1] have dominated the field, providing highly detailed simulations of network behavior. In contrast, ML-based models offer a scalable alternative to DES by leveraging data-driven approaches. GNNs have emerged as the dominant architecture in this space [19, 7], with the RouteNet family of models [5] leading the way. DES-ML hybrids aim to accelerate DES by replacing specific simulation components with ML-based approximations. Notable examples include MimicNet [3], DeepQueueNet [4] and m3 [20]. However, these models are typically trained on simulated data, which limits their accuracy and reliability in real-world networks without proper fine-tuning.

### v2

# Carlos Güemes-Palau, Miquel Ferriol-Galmés, Jordi Paillisse Vilanova, Albert López-Brescó, Pere Barlet-Ros and Albert Cabellos-Aparicio

@@ I Introduction
Network modeling is crucial for reliable communication, traffic optimization, and network design. Traditional approaches, such as Discrete Event Simulation (DES), have been the standard for accurately modeling network behavior. Tools like NS-3 [12] and OMNeT++ [15] simulate every network event. They provide highly detailed reconstructions of network dynamics but at a high computational cost, which limits scalability as networks grow. Recent advancements in Machine Learning (ML) have introduced transformative possibilities for network modeling, exemplified by models such as MimicNet [22], DeepQueueNet [17], and the RouteNet [ferriolgalmés2022routenetfermi] family. These innovations address the limitations of DES, enabling faster and more scalable network modeling at the price of a slight reduction in accuracy.
@@
Researchers often turn to network simulators to generate training datasets to mitigate data scarcity [22, 17, ferriolgalmés2022routenetfermi, 7]. Simulators offer a controlled and flexible environment for producing diverse and rare scenarios without the risks associated with live production networks. Conversely, generating simulated datasets is computationally expensive and often fails to capture the subtle nuances of real-world network dynamics. This arises from idealized assumptions in simulations and the proprietary nature of commercial network hardware. Consequently, ML models trained exclusively on simulated data exhibit reduced accuracy and reliability when deployed in production environments. This observation has been corroborated by other researchers [6] and by our findings in the evaluation (Table II).
This paper studies the viability of using simulated data to train ML models for real-world network behavior prediction. We propose a hybrid approach leveraging transfer learning [2, 1] to bridge the gap between simulation and real-world network environments. This consists in training and evaluating our approach by combining simulated network scenarios with real ones (Figure 1). We use RouteNet-Fermi [ferriolgalmés2022routenetfermi], a state-of-the-art ML network model, as our reference model in our comparison. However, we believe that our findings are generalizable to other ML-based solutions. We then use our configurable testbed network to record the real network scenarios. In our evaluation, we increase the accuracy of the resulting model by up to 88% relative to the version that only uses real-world network data. We also find that our approach is extremely data efficient: limiting the training process to 10 recorded network scenarios, using fine-tuning results in a model with $37\%$ less error than when training from scratch.
@@ II-A Transfer Learning
    Reused and adjustable weights: The weights are transferred from the donor model but are allowed to update during training and to better fit the target.
@@ II-B RouteNet-Fermi
RouteNet-Fermi [ferriolgalmés2022routenetfermi] is a state-of-the-art network performance model. Specifically, it uses a custom representation of the network along with a modified Message-Passing Neural Network architecture [5] that exploits the interactions between traffic flows and the underlying devices. As a result, it has proven to generate accurate predictions at a fraction of the computational cost of alternatives such as DES. Its design also makes it robust to infer under unseen topologies, including larger than those seen during the model’s training. It was also evaluated with both simulated and real-world network data.
@@ IV Methodology and experimental design
Our approach involves two main components: (1) training the RouteNet-Fermi [ferriolgalmés2022routenetfermi] model on simulated data and (2) fine-tuning it with real-world network data. This methodology enables us to combine the broad generalization capabilities gained from simulation with real-world specificity.
@@ IV-B Manual transfer learning
The resulting testbed configurations are listed in Table II. We split the network into blocks rather than individual layers to align with RouteNet-Fermi’s architecture. Unlike sequential NNs like MLPs, RouteNet-Fermi operates more like an ensemble of smaller NNs that work in parallel. For instance, the MLPs in the encoding block process individual network elements independently. Grouping these layers into blocks provides a structured approach to fine-tuning while ensuring that dependencies between blocks are respected. Furthermore, the shallow depth of the RouteNet-Fermi’s internal NNs, with the deepest component being a 3-layered MLP, limits the benefit of fine-tuning individual layers.
Figure 2: Visual example of fine-tuning a RouteNet-Fermi [ferriolgalmés2022routenetfermi] model, where the Encoding is frozen, the MPA is fine-tuned, and the Readout is re-trained.
@@ IV-C Automated transfer learning
  • Autofreeze [10]: This method begins by loading all donor weights and setting them as trainable. During training, blocks whose weight gradients fall under a threshold are frozen, minimizing computational costs.
@@
  • GTOT-Tuning [21]: A more advanced version of L2-SP meant for Graph Neural Networks (GNNs). Instead of comparing weights, it compares node embeddings after the MPA using the Masked Wasserstein Distance, which incorporates relational information.
@@ V Evaluation
TABLE I: Number of network scenarios available.
@@
Table I presents the number of network scenarios available for our study, categorized by source (simulated or real), traffic distribution (Poisson, On/Off, or MAWI), and partition (training, validation, or evaluation). Notably, the simulated dataset is significantly larger than the real-world dataset, reflecting the limited availability in real-world network data.
As part of the evaluation, we establish two baseline models: one trained exclusively on simulated scenarios and the other solely on real-world scenarios. We then use the simulated data only model for both manual and automated fine-tuning. Manual fine-tuning configurations and criteria for selecting these configurations are detailed in subsections IV-B. The automated fine-tuning approaches are listed in Section IV-C and have been implemented according to their original papers, including hyperparameter values, with necessary adaptations made to fit RouteNet-F’s architecture as needed. Our windowed implementation of the RouteNet-Fermi model is built upon their public repository [^1] https://github.com/BNN-UPC/RouteNet-Fermi, using TensorFlow 2.15. Model hyperparameters (e.g., embedding sizes) are borrowed directly from the original implementation. All models ran for a redundant number of epochs until the validation error stopped improving, and kept the weights that minimized it. Additional implementation details are available in our public repository [^2] https://github.com/BNN-UPC/Papers/wiki/Bridging-the-Gap-Between-Simulated-and-Real-Network-Data-Using-Transfer-Learning.
@@ V-A Fine-tuning the network model
Table II presents the MAPE for all evaluated fine-tuning configurations. Models are tasked to predict the average packet delay in each flow-window pair. The manual configurations involve varying the treatment of weights (to freeze, fine-tune, or re-train) in the three main blocks of the RouteNet-Fermi model: Encoding, MPA, and Readout. We also show the performance of automated finetuning methods and training exclusively on simulated data. The results are normalized against a model trained with the real-world network data only, without fine-tuning. That is, a value of 1 indicates the same accuracy as the baseline, while lower values indicate an improvement over the baseline.
@@ V-B Impact of real-world network data availability
The results show that with only five scenarios available, fine-tuning reduces the model’s MAPE from $19.01\%$ to $16.19\%$. Increasing the number of available scenarios to 10 further decreases the MAPE to $11.95\%$, a $37\%$ reduction. With 50 scenarios, the fine-tuned model achieves its largest improvement, lowering the MAPE to $9.95\%$, a $48\%$ decrease. This demonstrates the data efficiency enabled by fine-tuning. However, these results also reveal their limitations: beyond 125 scenarios, fine-tuning provides no additional benefit over training with real-world network data alone. While this threshold may vary between models, RouteNet-Fermi has demonstrated the ability to achieve high accuracy even with limited training data [ferriolgalmés2022routenetfermi]. This reinforces the idea that fine-tuning is particularly valuable when real-world network data is scarce.
@@ VI-A Network modeling
State-of-the-art network models can be broadly categorized into two fields: DES and ML-based models. DES simulators such as OMNeT++[15] and NS3[12] have dominated the field, providing highly detailed simulations of network behavior. In contrast, ML-based models offer a scalable alternative to DES by leveraging data-driven approaches. GNNs have emerged as the dominant architecture in this space [16, 7], with the RouteNet family of models [ferriolgalmés2022routenetfermi] leading the way. DES-ML hybrids aim to accelerate DES by replacing specific simulation components with ML-based approximations. Notable examples include MimicNet [22], DeepQueueNet [17] and m3 [8]. However, these models are typically trained on simulated data, limiting their accuracy and reliability in real-world networks.